OpenStudy (anonymous):

A company's monthly profit, P, from a product is given by P=-x^2+105x-1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1550?

OpenStudy (anonymous):

Please Help, This is what I have so far... 1550=-x^2+105x Am i on the right track? What do i do next?

OpenStudy (anonymous):

From what i understand you need to find the x when y = 1550. So you have an equation 1550 = -x^2+105x-1050 => x^2 - 105x +2600 = 0 Can you solve this?

OpenStudy (anonymous):

no im confused

OpenStudy (anonymous):

P=-x^2+105x-1050 this is a function. Let's use y instead of P y = -x^2 + 105x - 1050 x is the price in dollars and y is the resulting profit You need to find the lowest x to achieve $1550 profit. So set y = 1550 as i mentioned above and solve the quadratic equation.